OpenStudy (anonymous):

(a)The "lead" in pencils is a graphite composition with Young's modulus of about 1 x10^9 N/m^2. Calculate the change in length of the lead in an automatic pencil if you tap it straight into the pencil with a force of 4.0 N. The lead is 0.50 mm in diameter and 60 mm long.

OpenStudy (johnweldon1993):

Well remember how everything is connected \[\large \delta = E\epsilon \] And we know \[\large \delta = \frac{F}{A}\] and \[\large \epsilon = \frac{\Delta L}{L}\] So \[\large \frac{F}{A} = E\frac{\Delta L}{L}\] Just solve that for \(\large \Delta L\) \[\large \Delta L = \frac{FL}{AE}\] *You'll see later this is actually the deflection